OpenStudy (erinecookie):

Shannon manages a small zoo and she has been analyzing the attendance data. Shannon finds that the number of visitors increases exponentially as the temperature increases, and this situation is represented by the function f(x) = 3x. Shannon also finds a linear equation that models the number of people who leave the park early depending on the change in temperature, and it is represented by f(x) = −x + 4. The graph of the two functions is below. Find the solution to the two functions and explain what the solution represents.

OpenStudy (baru):

any ideas @erinecookie ?

OpenStudy (erinecookie):

I have no idea. Im so confused

OpenStudy (erinecookie):

I have no idea. Im so confused

OpenStudy (erinecookie):

Heres the graph with the question

OpenStudy (erinecookie):

I just need to find the solutions for F(x)=3^x and f(x)=-x+4 right? I dont know how to do it. But i know thats what i need to do

OpenStudy (baru):

exponentially is correct, you have typed your question wrong its not f(x)=3x it is f(x)=\(3^x\)

OpenStudy (erinecookie):

its f(x)=3^x my apologies

OpenStudy (baru):

ok, the solution to two functions is the point at which their graphs intersect. can you identify the point?

OpenStudy (erinecookie):

(1,3) ?

OpenStudy (baru):

yes :)

OpenStudy (erinecookie):

So thats the answer? The solution of both equations is (1,3)?

OpenStudy (baru):

yes that is the answer

OpenStudy (erinecookie):

Thank you so much!

OpenStudy (baru):

any idea what the solution means? x represents temperature f(x) represents number of people entering g(x) represents number of people leaving